Lars:

Which ordered pairs in the form (x, y) are solutions to the equation 4x−5y=24 Select each correct answer. A: (1, −4) B: (−9, −12) C: (6, 0) D: (4, 8)

Extrinix:

Simply put and very ironically, use the answer choices of the question to answer your question. You want to put all of these values into \(\sf{4x-5y=24}\) \(\sf{ A)~~4(1)-5(-4)~~~~~=24 \\ B)~~4(-9)-5(-12)=24 \\ C)~~4(6)-5(0)~~~~~~=24 \\ D)~~4(4)-5(8)~~~~~~~~=24 \\ }\) Then, you just solve for each of them, and if they equal up to \(\sf{24}\), then they are a valid point on the graph. I simplified the answer choices here for you: \(\sf{ A)~~~~~~~4~~+20~=24 \\ B)~~-36+60~=24 \\ C)~~~~~~~24+0~~~=24 \\ D)~~~~~~16~-40~=24 \\ }\) Best of luck, \(\style{font Family:Arial; text-shadow: 2px 2px 3px white, 0 0 10px white, 0 0 5px white;color:#23B7F3;font-size:100px;padding:50px;width:500px;height:100%;position:absolute;left:0;z-index:500;text-align:center;} {\Huge\sf\color {} {{-~Akuma~matata}}}\) "what a wonderful phrase"

Lars:

Thank you, that makes sense.

Extrinix:

No problem! ^^